# Emergent Dominance Hierarchies in Reinforcement Learning Agents

### 27 Feb 2024 v4

@@ 1. Introduction
First, we adapt the ethological theory of dominance hierarchies to artificial agents by modeling conflicts between each pair of individuals as the classic game of Chicken (Rapoport and Chammah, 1966). [^1] We use the word ”chicken” in this paper to refer to both the Game Theory concept and the beak-and-feathers animal. We capitalize ”Chicken” to indicate the former. To accommodate groups of more than two individuals, we generalize the game of Chicken into an $N$-player stochastic game that we call Chicken Coop. We borrow the established terminology and definitions from animal study to allow calculating dominance metrics from agents’ action history. We release the Chicken Coop environment under the MIT open-source license. [^2] The code and installation instructions are available at https://github.com/cool-RR/chicken-coop
@@ 3.2. Applying dominance hierarchies to MAS/MARL
Vezhnevets et al. (2023) and Wu et al. (2023) replace the RL algorithms used by the agents with large language models (LLMs), which have gained immense popularity due to their robustness in solving a wide variety of problems (Xi et al., 2023). These LLM-based agents are organized in human-inspired hierarchies such as manager-workers and student-assistant-expert. The back-and-forth interaction between multiple LLM agents is effective at encouraging diversity (Liang et al., 2023), factuality, and reasoning capability (Du et al., 2023), when compared to the use of a single LLM-based agent. The potential usefulness of such agents is tremendous, as LLMs can be seamlessly integrated with multimodal foundation models (Li et al., 2023), allowing them to process and generate textual, visual and auditory content. We suggest that a formal understanding and terminology of agent hierarchies could play a decisive role in maximizing the effectiveness of these agents.
@@ 4.1. Dominance between two agents
Agent 1 Agent 2 Hawk Dove
Hawk $C,C$ $W,L$
Dove $L,W$ $T,T$
The reward constants satisfy $W>T>L>C$. These are short for win, tie, lose and crash. The $(L,W)$ and $(W,L)$ outcomes serve as the Nash equilibria $NE_{i}$ and $NE_{j}$ Each agent’s aggressiveness is simply its tendency to play hawk.
@@ 4.2. Dominance hierarchies
Chicken Coop is an $N$-player generalization of the game of Chicken. In each episode of Chicken Coop, the agents are divided into random pairs that are used for the entire episode. Each pair of agents plays one round of Chicken against each other, choosing either hawk or dove and receiving a reward in $\{W,T,L,C\}$. Each agent’s sole observation is the identity of their opponent agent.
@@ 5. Methods
In all of the experiments below, unless stated otherwise, the RL agents are trained using the PPO algorithm (Schulman et al., 2017) implementation provided by the RLlib framework (Liang et al., 2018). We train $L=300$ populations of $N=6$ Chicken Coop agents, at a learning rate of $2\times 10^{-6}$, a discount factor of $\gamma=0.99$, and a clipping parameter of $\epsilon=0.3$.
@@ 6.3. Observation ablation
Methodology: We define an agent’s opponent perception accuracy (OPA) as a number between 00 and $1$. For each agent in each episode, the probability that its observation of its opponent is replaced with a random index number is $(1-OPA)$. For each OPA value in $\{0,0.1,0.2,\ldots,1\}$ we train $L=10$ different population of $N=6$ Chicken Coop agents at a learning rate of $2\times 10^{-6}$ and measure their rapport.
@@ 6.4. Transmitting dominance hierarchies to new populations
    Each agent acts as an independent enforcer of its dominance hierarchy, punishing any other agents that disobey it.
@@ 7. Discussion and future work
    Opponent Shaping: Algorithms such as LOLA (Foerster et al., 2018) and M-FOS (Lu et al., 2022) enable agents to take the learning process of the other agents into account in their own policy gradients. When applied to an environment that supports dominance hierarchies, they may promote second-order dominance-seeking strategies, e.g., they may prompt agent $i$ to consider how it can behave as to encourage agent $j$ to place agent $i$ at a high rank in the dominance hierarchy.
When humans work on problems as a group, we balance in-group intrigues with external pressures to provide good results. We suggest that this interplay between individual needs and group needs may play a crucial role in the success of our collective intelligence. Therefore, we propose augmenting a population of dominance-seeking agents with multimodal foundation models (Li et al., 2023), and giving them real-world tasks to work on collaboratively. We hypothesize that the real-world decisions made by such agents may be more interpretable and corrigible, as human operators may recognize that the agents’ decision process reflects their own.

### 21 Mar 2024 v5

@@ 1. Introduction
First, we adapt the ethological theory of dominance hierarchies to artificial agents by modeling conflicts between each pair of individuals as the classic game of Chicken (Rapoport and Chammah, 1966). [^1] We use the word “chicken” in this paper to refer to both the Game Theory concept and the beak-and-feathers animal. We capitalize “Chicken” to indicate the former. To accommodate groups of more than two individuals, we generalize the game of Chicken into an $N$-player stochastic game that we call Chicken Coop. We borrow the established terminology and definitions from animal study to allow calculating dominance metrics from agents’ action history. We release the Chicken Coop environment under the MIT open-source license. [^2] The code and installation instructions are available at https://github.com/cool-RR/chicken-coop
@@ 4.1. Dominance between two agents
Agent 1 Agent 2 Dove Hawk
Dove $R,R$ $S,T$
Hawk $T,S$ $P,P$
The reward constants satisfy $T>R>S>P$. The $(S,T)$ and $(T,S)$ outcomes serve as the Nash equilibria $NE_{i}$ and $NE_{j}$ Each agent’s aggressiveness is simply its tendency to play hawk.
@@ 4.2. Dominance hierarchies
Chicken Coop is an $N$-player generalization of the game of Chicken. In each episode of Chicken Coop, the agents are divided into random pairs that are used for the entire episode. Each pair of agents plays one round of Chicken against each other, choosing either hawk or dove and receiving a reward in $\{T,R,S,P\}$. Each agent’s sole observation is the identity of their opponent agent.
@@ 5. Methods
In all of the experiments below, unless stated otherwise, the RL agents are trained using the PPO algorithm (Schulman et al., 2017) implementation provided by the RLlib framework (Liang et al., 2018). We train $L=300$ populations of $N=6$ Chicken Coop agents, at a learning rate of $2\times 10^{-6}$, a discount factor of $\gamma=0.99$, and a clipping parameter of $\epsilon=0.3$. We use a dominance threshold of $\eta=0.55$ and these reward constants:
$R=0,\quad S=-2,\quad T=5,\quad P=-10$
@@ 6.3. Observation ablation
Methodology: We define an agent’s opponent perception accuracy (OPA) as a number between 00 and $1$. For each agent in each episode, the probability that its observation of its opponent is replaced with a random index number is $(1-OPA)$. For each OPA value in $\{0,0.1,0.2,\ldots,1\}$ we train $L=10$ different populations of $N=6$ Chicken Coop agents at a learning rate of $2\times 10^{-6}$ and measure their rapport.
@@ 6.4. Transmitting dominance hierarchies to new populations
    Each agent learns to consistently play $hawk$ with all its subordinates, and this tendency serves as a punishing mechanism (Axelrod, 1986), or a sanction. Any errant attempt by a subordinate to play $hawk$ against a dominant would usually result in a $hawk/hawk$ outcome, with both agents receiving the lowest possible reward of $P$. All non-bottom agents effectively act as independent enforcers of their dominance hierarchy, punishing any agent that disobeys it.
@@ 7. Discussion and future work
    Opponent Shaping (OS): Algorithms such as LOLA (Foerster et al., 2018) and M-FOS (Lu et al., 2022) enable agents to take the learning process of the other agents into account in their own policy gradients. When applied to an environment that supports dominance hierarchies, they may promote second-order dominance-seeking strategies, e.g., they may prompt agent $i$ to consider how it can behave as to encourage agent $j$ to place agent $i$ at a high rank in the dominance hierarchy.
We propose that the study of dominance hierarchies between RL agents may be combined with research on large language models (LLMs) and multimodal foundation models (MFMs), which have gained immense popularity due to their robustness in solving a wide variety of problems (Xi et al., 2023; Li et al., 2023). Tufano et al. (2024) and similar work (Vezhnevets et al., 2023; Wu et al., 2023) present an innovative way to use LLMs: instead of sending queries to a single LLM, users converse with a group of LLM-based agents, organized in human-inspired hierarchies such as manager-workers and student-assistant-expert. The back-and-forth interaction between multiple LLM agents is effective at encouraging diversity (Liang et al., 2023), factuality, and reasoning capability (Du et al., 2023), when compared to a single LLM-based agent. We suggest that the augmenting these LLM-based agents with RL or OS algorithms, specifically to maintain the social relationships between the agents, may result in hierarchies that are similar to those observed in biological life.
When humans work on problems as a group, we balance in-group intrigues with external pressures to provide good results. We suggest that this interplay between individual needs and group needs may play a crucial role in the success of our collective intelligence. We hypothesize that MFM-based agents operating under similar conditions may make real-world decisions in a way that is more interpretable and corrigible, as human operators may recognize that the agents’ decision process reflects their own.